OpenStudy (anonymous):

Two lines, A and B, are represented by equations given below: Line A: y = x – 4 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? A. (–3, –5), because the point satisfies one of the equations B. (–3, –5), because the point lies between the two axes C. (–4, –8), because the point satisfies both equations D. (–4, –8), because the point does not lie on any axis

OpenStudy (anonymous):

You want the value of y and x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Line A : y = x - 4 - eqn1 Line B : y = 3x + 4 - eqn2 Plug eqn 1 in eqn 2 x - 4 = 3x + 4 So far , good?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

then how do i know witch one is the answer?

OpenStudy (anonymous):

x - 4 = 3x + 4 x - 3x = 4 + 4 -2x = 8 x =8/-2 x = -4

OpenStudy (anonymous):

Plug x = -4 in either eqn 1 or eqn 2 y = x - 4 y = -4 - 4 y = -8

OpenStudy (anonymous):

because the point satisfies both equations?

OpenStudy (anonymous):

or

OpenStudy (anonymous):

because the point does not lie on any axis?

OpenStudy (loser66):

hey, why do you need more help? you got (-4,-8) and there is ONLY one option looks like this. What is next?? PICK IT. dat sit

OpenStudy (anonymous):

but i don't know wich on it is: because the point does not lie on any axis? or because the point satisfies both equations?

Vocaloid (vocaloid):

a solution to a system of equations is a point that satisfies both equations

OpenStudy (loser66):

YES, LIKE wshat Vocaloid said